OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

x2 + 6x + 8

OpenStudy (anonymous):

how it would be solved by graphing, factoring, the quadratic formula, and by completing the square.

OpenStudy (anonymous):

okay completeing teh square: x2+6x +8=0 x2 +6x = -8 move the 8 over use this formula (b/2)~2 so its (6/2)~2 oe (3)~2 = 9 x2 +6x +9 = -8 +9 add the 9 to both sides (x+3)~2 = 1 \[(x+3)^{2} = 1\]

OpenStudy (anonymous):

then

OpenStudy (anonymous):

\[\sqrt{(x+3)^{2}} = \pm \sqrt{1}\] sqaure root both sides

OpenStudy (anonymous):

x+3 = + or - 1 -3 -3 subtract 3 from both sides x=+1 -3 and -1-3 or -2, -4

OpenStudy (anonymous):

and thats the answer

OpenStudy (anonymous):

x= -2, -4 yes

OpenStudy (anonymous):

for completing the sqaure

OpenStudy (anonymous):

thank you you've been a big help

OpenStudy (anonymous):

great